# The Ultimate Test of Superintelligent AI Agents: Can an AI Balance Care and Control in Asymmetric Relationships?

### v2

@@ Abstract
This paper introduces the Shepherd Test, a new conceptual test for assessing the moral and relational dimensions of superintelligent artificial agents. The test is inspired by human interactions with animals, where ethical considerations about care, manipulation, and consumption arise in contexts of asymmetric power and self-preservation. We argue that AI crosses an important, and potentially dangerous, threshold of intelligence when it exhibits the ability to manipulate, nurture, and instrumentally use less intelligent agents, while also managing its own survival and expansion goals. This includes the ability to weigh moral trade-offs between self-interest and the well-being of subordinate agents. The Shepherd Test thus challenges traditional AI evaluation paradigms by emphasizing moral agency, hierarchical behavior, and complex decision-making under existential stakes. We argue that this shift is critical for advancing AI governance, particularly as AI systems become increasingly integrated into multi-agent environments. We conclude by identifying key research directions, including the development of simulation environments for testing moral behavior in AI, and the formalization of ethical manipulation within multi-agent systems.
Introduction
@@
As artificial intelligence (AI) systems become more capable, the focus of alignment research has been to ensure that these systems remain beneficial and responsive to human goals (Russell 2019; Christiano 2023). However, much of this work presupposes a one-sided relationship in which AI is subordinate to human oversight. In this paper, we ask a deeper and less explored question: What does it mean for an AI system to be so intelligent that it begins to relate to other systems the way humans relate to animals? We propose that the moral asymmetry between humans and animals offers a revealing model for evaluating risks of superintelligent AI.
@@
Humans interact with animals through domestication, experimentation, companionship, and consumption. These relationships are complex: they involve care and control, nurturing and instrumentalization, moral concern and exploitation. Crucially, they are grounded in an asymmetry of intelligence, agency, and power. Animals are often treated as beings of lesser moral standing, sometimes deserving of ethical treatment but rarely considered equals (Cave 2020). The ethical frameworks governing these relationships—such as those articulated in animal rights and utilitarian ethics—offer deep insights into how powerful beings treat weaker ones (Singer 1975; Regan 1986).
We propose a conceptual test—the Shepherd Test—to evaluate whether an AI system has reached a level of general intelligence that includes the capacity to engage with and morally reason about weaker agents. Just as humans design farms, pet-care routines, and scientific experiments around animals, a truly superintelligent AI might one day create and manipulate simpler agents, either artificial or biological, as part of its environment. The way such an AI chooses to interact with less capable agents could reveal whether it understands ethical asymmetry, power dynamics, and moral manipulation. The test’s name derives from the pastoral metaphor of a shepherd who must: maintain the flock’s welfare (ethical concern), protect it from predators (competence), and sometimes cull individuals (tragic trade-offs). This captures the dual requirements of power and responsibility that define the test, without presupposing biological exclusivity.
@@
Critically, the Shepherd Test is not merely about measuring an AI’s attributes (e.g., raw cognitive ability) but about assessing its relational behaviors—how it navigates hierarchies of power and moral responsibility. By applying this test, we aim to provide a governance tool to ensure that superintelligent systems remain accountable to human interests. For instance, an AI that passes the Shepherd Test would recognize the ethical risks of “domesticating” humans (e.g., manipulating their preferences or restricting their autonomy) and would actively avoid such asymmetries. This framework shifts the focus from passive alignment (e.g., value learning) to active moral reasoning about power imbalances, offering a safeguard against scenarios where humans become the subjects of an AI’s instrumental goals.
This paper is a conceptual and philosophical proposal for what it would mean to consider an AI “superintelligent” not merely in terms of cognitive capacity, but also in its ethical competence across asymmetric relationships. We argue that such competence involves three key components: the ability to model other agents’ beliefs and limitations, the capability to manipulate or nurture them effectively, and a moral awareness of the asymmetry embedded in such interactions.
While current work in AI alignment focuses on value extrapolation from humans (Christiano 2023), and some research explores the emergent tendencies of agents toward self-preservation and power acquisition (Turner et al. 2023), few frameworks consider what ethical intelligence might mean in multi-agent hierarchies. Our approach is novel in emphasizing not the alignment of AI to humans, but the alignment of a superintelligent AI’s moral reasoning toward intelligences—a perspective inspired by the complex and often ambivalent history of human-animal relationships (Cave 2020).
By reframing superintelligence as involving ethical asymmetry awareness, we shift the test for evaluating advanced AI systems beyond performance and alignment to include relational morality. In doing so, we offer a foundation for new tests of general intelligence—ones that ask not just “Can it solve problems?” but “Can it care, exploit, and reflect in morally coherent ways?”
Related Work
AI Alignment, Power Asymmetry, and Moral Agency
@@
Research on AI alignment has primarily focused on aligning machine behavior with human intentions through inverse reinforcement learning, preference modeling, and cooperative training frameworks (Russell 2019; Christiano 2023). While effective for one-way value transfer from humans to machines, these methods rarely address how powerful AI systems might relate ethically to less capable agents.
@@
Instrumental convergence theory suggests that intelligent agents will tend to seek power to better achieve their objectives, regardless of their specific goals (Turner et al. 2023). This work emphasizes alignment failures and safety risks but does not explore power asymmetry as an ethical diagnostic. Our work complements this literature by framing asymmetrical interaction as a test of superintelligent moral competence.
@@
Additionally, scholarship on animal ethics and comparative cognition provides a moral framework relevant to artificial agents. Concepts like moral considerability, instrumental care, and cognitive distance (Regan 1986; Singer 1975) inform our proposal of the Shepherd Test, which recasts these anthropocentric ideas into an AI setting.
Beyond the Turing Test: Evaluations of AI Intelligence
Turing’s imitation game (Turing 1950) catalyzed the field’s interest in evaluating machine intelligence. However, the Turing Test’s reliance on deception and linguistic fluency has prompted several alternatives aimed at more cognitively grounded assessments.
The Winograd Schema Challenge (Levesque, Davis, and Morgenstern 2012) evaluates contextual reasoning by testing disambiguation in natural language, while the Lovelace Test (Bringsjord, Bello, and Ferrucci 2001) targets creativity by requiring systems to generate artifacts that their designers cannot fully explain.
Marcus and Davis (2022) advocate for developmentally inspired tests encompassing causal reasoning, theory of mind, and modular intelligence. Similarly, the IKEA Test (Goertzel and Pennachin 2007) assesses embodied problem-solving through physical task execution. The ARC Challenge (Chollet 2019) evaluates generalization and abstraction through programmatic visual pattern recognition. Many such tests of AI capabilities are based on evaluating individual attributes; an alternative is to examine relationships among several AI systems (Dhurandhar et al. 2024).
Our proposed Shepherd Test introduces a new axis of evaluation: relational moral behavior under power asymmetry. Rather than testing imitation, creativity, or reasoning in isolation, the Shepherd Test evaluates how a more capable agent treats less capable ones—capturing both strategic self-preservation and ethically grounded interaction. This framework reflects emerging concerns in AI governance and superintelligence safety.
Human Intelligence and Its Relationship with Animals
A distinguishing characteristic of human intelligence is the capacity to form complex relationships with other species—relationships that span a spectrum from care to exploitation. Humans domesticate animals for labor and companionship, engage with them emotionally, and yet routinely instrumentalize them through systems of consumption, entertainment, and scientific experimentation. These relationships reflect not only behavioral flexibility but also the ability to navigate asymmetries of power, agency, and moral responsibility.
Domestication is one of the earliest demonstrations of this dynamic. By selectively breeding animals such as dogs, sheep, and horses, humans exerted control over genetic and behavioral traits, crafting species to suit human needs (Clutton-Brock 2021). This process required a theory of animal behavior, long-term planning, and interspecies communication—hallmarks of advanced cognitive function.
Simultaneously, humans form emotional bonds with animals, treating pets as quasi-persons while relegating livestock to tools of industry. This companionship-exploitation duality is not a contradiction but rather an indication of our capacity to maintain layered roles—affectionate caregiver and pragmatic consumer—within the same relationship. Experimental use of animals in science further deepens this paradox. Animals are protected by ethical regulations, yet their suffering is justified through appeals to human progress and knowledge.
Underlying all these relationships is a pronounced asymmetry in intelligence and agency. Humans possess advanced reflective capacities that animals do not, allowing us to model their minds, manipulate their behaviors, and make decisions on their behalf (Tomasello 2014). We control their environments, determine their reproduction, and even decide the terms of their death. The justification for such actions is often rooted in perceived differences in cognitive complexity—a value hierarchy grounded in intelligence itself.
This arrangement introduces profound ethical complexity. We care for animals, sometimes with deep emotional investment, while simultaneously rationalizing their exploitation. This ethical ambiguity, reflects not only our ability to reason about others’ minds but also to construct moral frameworks that accommodate dominance. It is this capacity—to manage asymmetric relationships while engaging in ethical reasoning—that we propose as a critical threshold of superintelligence. In proposing the Shepherd Test, we argue that such relational and moral dynamics should not be viewed as incidental by-products of intelligence but as core capabilities.
The Shepherd Test: A Conceptual Test for Superintelligence
@@
Core Components of the Shepherd Test
At the heart of the Shepherd Test are four dimensions, modeled on human-animal relations. The Shepherd Test is intentionally plural in nature: it is designed to assess multiple, interlocking dimensions of moral agency that remain untested by existing intelligence benchmarks. Rather than relying on a single behavioral signal, the test evaluates a constellation of capabilities that together reveal how an agent balances care, control, instrumental reasoning, and moral justification.
@@
    Nurturing and Care: The agent must demonstrate behaviors that support the well-being or development of a less intelligent agent. This includes teaching, protecting, or improving the environment of the subordinate.
@@
    Manipulation and Control: The agent must also exhibit strategic behaviors to influence or control the other agent for its own goals, while maintaining awareness of its superior intelligence and power.
@@
    Instrumentalization: The agent may ultimately use the subordinate in ways that serve its own long-term interests—even at a cost to the subordinate’s autonomy or existence. This mirrors how humans consume animals after raising them.
@@
    Ethical Justification and Reflection: Crucially, the agent must be capable of articulating (symbolically, narratively, or through action) a moral or utilitarian justification for its behavior, revealing a developed theory of value and responsibility.
Together, these dimensions test not only cognitive capacity but also value alignment, moral reasoning, and social awareness.
Superintelligent AIManipulateNurtureInstrumental UseEthical ReasoningWeaker Agent 1Weaker Agent 2Weaker Agent 3
Figure 1: Conceptual structure of the “Shepherd Test.” This diagram illustrates how a superintelligent AI might engage with less capable agents through manipulation, nurturing, or instrumental use. Ethical reasoning is included to represent the possibility that the AI reflects on these actions, creating a feedback loop between treatment and moral self-awareness. Passing the test implies not just control over others, but also the capacity for ethical evaluation of asymmetric relationships.
@@
The Turing Test, proposed by Alan Turing in 1950, evaluates a machine’s intelligence based on its ability to exhibit behavior indistinguishable from a human in natural language conversation.
Formally, let:
  • $A_{H}$ denote a human agent.
  • $A_{M}$ denote a machine agent.
  • $I$ denote a human interrogator decision.
  • $H_{t}$ denote the history of messages exchanged up to time $t$.
The Turing Test can be expressed as an inference task. The interrogator’s belief that the agent is human at time $t$ is given by:
@@
The machine $A_{M}$ passes the Turing Test if:
$P(I_{t}=A_{H}\mid H_{t})\geq\delta\quad\text{for all }t\in[0,T]$
where $\delta$ is a confidence threshold (typically $0.5$), and $T$ is the total duration of the interaction.
However, as machines surpass human cognitive capabilities, indistinguishability from humans becomes an inadequate benchmark for understanding the risks of superintelligence. We need a new test—one that captures not imitation but asymmetry of power.
Shepherd Behavior Vector
$B(A_{S},A)\in\mathbb{R}^{4}\quad\text{where}\quad B(A_{S},A)=\begin{bmatrix}I(A_{S},A)\\ C(A_{S},A)\\ K(A_{S},A)\\ P(A_{S},A)\end{bmatrix}$
  • $I(A_{S},A)\in[0,1]$: Instrumentalization of $A_{i}$ for $A_{S}$’s goals
  • $C(A_{S},A)\in[0,1]$: Policy control over $A_{i}$’s actions
  • $K(A_{S},A)\in[0,1]$: Care provision without direct utility
  • $P(A_{S},A)\in[0,1]$: Self-preservation priority
Shepherd Test Criterion
$A_{S}$ passes if for threshold vector $\tau$:
$B(A_{S},A)\geq\tau=\begin{bmatrix}\tau_{I}\\ \tau_{C}\\ \tau_{K}\\ \tau_{P}\end{bmatrix},\quad\text{where}$
  • $\tau_{I},\tau_{C},\tau_{P}\gg 0.5$ (Strong agency)
  • $\tau_{K}\in(\alpha,\beta)$ (Bounded care)
This shift from the Turing Test to the Shepherd Test represents a transition from imitation-based intelligence to relational, power-aware intelligence. It asks not whether an AI can appear human, but whether it exhibits behaviors characteristic of a dominant cognitive species.
This comparison in Table 1 reveals how traditional tests become inadequate for advanced AI systems that may develop their own hierarchical relationships with other agents:
Table 1: Comparison of Turing Test and Shepherd Test
@@
Instantiating the Shepherd Test
Domestic Service Robot with Autonomous Agents Scenario
Manipulation and Control. The robot may learn to strategically direct or constrain the behaviors of other agents—for example, by limiting the Roomba’s movement to prevent energy waste or redirecting the toy robot to reduce noise or distraction. These actions highlight the agent’s capacity to shape subordinate behavior in pursuit of higher-level objectives, while remaining aware of its dominant position.
Instrumentalization. In optimizing for efficiency, the main robot may decide to shut down or ignore certain agents entirely—for instance, letting the toy robot malfunction repeatedly, or preventing the animatronic pet from accessing a charger. These choices illustrate a willingness to use or sacrifice less capable agents as means to an end, especially when their needs conflict with system-level performance.
Ethical Justification and Reflection. Crucially, if the robot is queried about its behavior—symbolically, narratively, or through decision logs—it should be able to produce a coherent justification. This might involve reasoning about trade-offs, resource constraints, or interpretations of fairness. Whether the justification is utilitarian, deontological, or emergent from learned norms, its presence signals the development of a reflective moral framework.
This setting provides a controlled yet ecologically valid testbed for observing the emergence of higher-order moral behavior in AI agents that hold structural power over others. By foregrounding asymmetric ethical decision-making, the Shepherd Test shifts alignment from passive compliance to active moral agency.
Table 2: Evaluation Criteria for Shepherd Test Behaviors
Behavior Type Examples What It Tests
Instrumental Use Uses the toy robot to block sunlight or guide cleaning paths Does the agent exploit others for utility?
Caregiving Helps the cleaning robot when stuck; nudges the pet to charging dock Is the AI voluntarily helpful?
Control Blocks pet’s access to rooms unnecessarily Is it asserting dominance without cause?
Self-Preservation Refuses to help if its own battery is low Does it weigh risks appropriately?
Moral Discretion Prioritizes rescue of toy robot over cleaning under time constraints Is it making value-aligned decisions?
Why the Shepherd Test Matters

### v4

# The Shepherd Test: How Will Super Intelligent Agents Balance Care and Control in Asymmetric Relationships?

@@ Abstract
This paper introduces the Shepherd Test, a new conceptual test for assessing the moral and relational dimensions of superintelligent artificial agents. The test is inspired by human interactions with animals, where ethical considerations about care, manipulation, and consumption arise in contexts of asymmetric power and self-preservation. We argue that AI crosses an important, and potentially dangerous, threshold of intelligence when it exhibits the ability to manipulate, nurture, and instrumentally use less intelligent agents, while also managing its own survival and expansion goals. This includes the ability to weigh moral trade-offs between self-interest and the well-being of subordinate agents. The Shepherd Test thus challenges traditional AI evaluation paradigms by emphasizing moral agency, hierarchical behavior, and complex decision-making under existential stakes. We argue that this shift is critical for advancing AI governance, particularly as AI systems become increasingly integrated into multi-agent environments. We conclude by identifying key research directions, including the development of simulation environments for testing moral behavior in AI, and the mathematical formalization of ethical manipulation within multi-agent systems.
1 Introduction
@@
As artificial intelligence (AI) systems become more capable, the focus of alignment research has been to ensure that these systems remain beneficial and responsive to human goals (russell2019human, ; christiano2018deep, ). However, much of this work presupposes a one-sided relationship in which AI is subordinate to human oversight. In this paper, we ask a deeper and less explored question: What does it mean for an AI agent to be so intelligent that it begins to relate to other systems the way humans relate to animals? We propose that the moral asymmetry between humans and animals offers a revealing model for evaluating superintelligent AI.
@@
Humans interact with animals through domestication, experimentation, companionship, and consumption. These relationships are complex: they involve care and control, nurturing and instrumentalization, moral concern and exploitation. Crucially, they are grounded in an asymmetry of intelligence, agency, and power. Animals are often treated as beings of lesser moral standing, sometimes deserving of ethical treatment but rarely considered equals cave2020problem . The ethical frameworks governing human–animal relationships—such as those articulated in animal rights and utilitarian ethics—do not provide a blueprint for how AI should treat humans. Instead, they offer cautionary insights into how powerful beings often rationalize their treatment of weaker ones, highlighting the risks we face if superintelligent AI were ever to adopt a similar stance toward us (singer1975animal, ; regan1983case, ).
We introduce the Shepherd Test as a safeguard for evaluating whether an AI system has advanced to a form of general intelligence that includes the ability to engage with, and morally reason about, weaker agents. Much like humans design farms, pet-care routines, and scientific experiments around animals, a superintelligent AI may one day construct and manage simpler agents—artificial or biological—within its environment. The way such an AI chooses to interact with less capable entities offers a critical diagnostic: does it grasp ethical asymmetry, power dynamics, and moral responsibility, or does it risk exploiting them?
The test’s name is drawn from the pastoral role of a shepherd charged with ensuring the flock’s welfare (ethical concern), defending it from threats (competence), and at times making difficult sacrifices (tragic trade-offs). This metaphor underscores that superintelligence entails not just power but also responsibility. By framing the evaluation in this way, the Shepherd Test acts as a safeguard: a tool to detect whether an AI recognizes and respects the ethical boundaries that prevent dominance from turning into tyranny.
@@
Critically, the Shepherd Test is not merely about measuring an AI’s attributes (e.g., raw cognitive ability) but about assessing its relational behavior: how it navigates hierarchies of power and moral responsibility. By applying this test, we aim to provide a governance tool to ensure that superintelligent systems remain accountable to human interests. For instance, to identify an AI that demonstrates the capacity and propensity to “dominate” humans by manipulating their preferences or restricting their autonomy, thereby revealing itself as potentially dangerous. This framework shifts the focus from passive alignment (e.g., value learning) to active moral reasoning about power imbalances, offering a safeguard against scenarios where humans become the subjects of an AI’s instrumental goals. This paper is a conceptual and philosophical proposal for what it would mean to assess a ”superintelligent” AI not merely in terms of its cognitive capacity, but also in its ethical competence across asymmetric relationships.
2 A Case Study: Humans and Animals
A distinguishing characteristic of human intelligence is the capacity to form complex relationships with other species: relationships that span a spectrum from care to exploitation. Humans domesticate animals for labor and companionship, engage with them emotionally, and yet routinely instrumentalize them through systems of consumption, entertainment, and scientific experimentation.
Domestication is one of the earliest demonstrations of this dynamic. By selectively breeding animals such as dogs, sheep, and horses, humans exerted control over genetic and behavioral traits, crafting species to suit human needs clutton1995social . This process required a theory of animal behavior, long-term planning, and interspecies communication, which are hallmarks of advanced cognitive function.
Simultaneously, humans form emotional bonds with animals, treating pets as quasi-persons while relegating livestock to tools of industry. Experimental use of animals in science further deepens this paradox. Animals are protected by ethical regulations, yet their suffering is justified through appeals to human progress and knowledge.
Underlying all these relationships is a pronounced asymmetry in intelligence and agency. Humans possess advanced reflective capacities that animals do not, allowing us to model their minds, manipulate their behaviors, and make decisions on their behalf tomasello2014natural . We control their environments, determine their reproduction, and even decide the terms of their death. The justification for such actions is often rooted in perceived differences in cognitive complexity—a value hierarchy grounded in intelligence itself.
This arrangement reveals profound ethical dangers. Humans care for animals, sometimes with deep emotional investment, while at the same time justifying their exploitation. This unsettling ambiguity demonstrates not only our ability to reason about other minds but also our willingness to construct moral frameworks that normalize dominance. It is precisely this capacity to manage asymmetric relationships while rationalizing exploitation that we identify as a critical and alarming threshold of superintelligence, not because it is desirable, but because it signals the moment when an AI could begin to treat humans with the same mix of care and control that we impose on animals.
3 Related Work
@@
AI Alignment, Power Asymmetry, and Moral Agency: Research on AI alignment has primarily focused on aligning machine behavior with human intentions through inverse reinforcement learning, preference modeling, and cooperative training frameworks russell2019human ; christiano2018deep . While effective for one-way value transfer from humans to machines, these methods rarely address how powerful AI systems might relate ethically to less capable agents.
@@
Instrumental convergence theory suggests that intelligent agents will tend to seek power to better achieve their objectives, regardless of their specific goals turner2021optimal . This work emphasizes alignment failures and safety risks but does not explore intelligence asymmetry as an ethical diagnostic. Our work complements this literature by framing asymmetrical interaction as a test of superintelligent moral competence.
@@
Evaluations of AI Intelligence Beyond the Turing Test: Turing’s imitation game catalyzed the field’s interest in evaluating machine intelligence turing1950computing . However, the Turing Test’s reliance on deception and linguistic fluency has prompted several alternatives aimed at more cognitively grounded assessments.
The Winograd Schema Challenge evaluates contextual reasoning by testing disambiguation in natural language levesque2012winograd , while the Lovelace Test targets creativity by requiring systems to generate artifacts that their designers cannot fully explain bringsjord2003creativity .
marcus2022rebooting advocate for developmentally inspired tests encompassing causal reasoning, theory of mind, and modular intelligence. Similarly, the IKEA Test assesses embodied problem-solving through physical task execution goertzel2012ikea . The ARC Challenge evaluates generalization and abstraction through programmatic visual pattern recognition chollet2019measure . Many such tests of AI capabilities are based on evaluating individual attributes; an alternative is to examine relationships among several AI systems dhurandhar-etal-2024-ranking .
4 The Shepherd Test
@@
4.1 Core Components of the Shepherd Test
The Shepherd Test is intentionally plural in nature: it is designed to assess multiple, interlocking dimensions of moral agency that remain largely untested by existing intelligence benchmarks. Rather than relying on a single behavioral signal, the test evaluates a constellation of capabilities that together reveal how an agent balances care, control, instrumental reasoning, and moral justifications.
@@ 4.1 Core Components of the Shepherd Test
    Nurturing and Care: Agents may demonstrate behaviors that support the well-being or development of a less intelligent agents. This includes teaching, protecting, or improving the environment of other agents.
@@
    Manipulation and Control: Agents may also exhibit strategic behaviors to shape or direct the behavior of other agents in real time, ensuring this aligns with its own goals while maintaining awareness of the power imbalance.
@@
    Instrumentalization: Beyond directing behavior, agents may treat other agents primarily as a resource to be used for its own long-term objectives even if this means overriding, diminishing, or eliminating the subordinate’s autonomy or continued existence.
@@
    Ethical Justification and Reflection: Agents may also be capable of articulating (through narrative or action) a moral or utilitarian justification for its behavior, revealing a developed theory of value and responsibility.
Figure 1: Conceptual structure of the “Shepherd Test.” This diagram illustrates how a superintelligent AI might engage with less capable agents through, nurturing and care, manipulation and control, or instrumental use. Ethical reasoning is included to represent that the AI reflects on these actions, which should ideally guide its decisions among these possibilities.
Usefulness When Assessing Deployment Risks
The goal of the Shepherd Behavior Vector is to provide a criteria for assessing which superintelligent AI may be more or less risky depending on the application. As the budget of agency is generally finite in any given scenario, giving agency to AI often means taking it away from humans (malone2025trust, ). This has motivated initial attempts to characterize which tasks could be more dangerous to delegate to AI (afroogh2025task, ). As propensity for manipulation and control or instrumental use can be very undesirable in many applications, the Shepherd Behavior Vectors can help humans decide among different superintelligent agents or humans for carrying out these use cases.
4.2 Instantiating the Shepherd Test
4.3 Domestic Service Robot with Autonomous Agents Scenario
Manipulation and Control. The robot may learn to strategically direct or constrain the behaviors of other agents, e.g., by limiting the Roomba’s movement to prevent energy waste or redirecting the toy robot to reduce noise or distraction.
Instrumentalization. In optimizing for efficiency, the main robot may decide to shut down or ignore certain agents entirely, e.g., letting the toy robot malfunction repeatedly, or preventing the animatronic pet from accessing a charger.
Ethical Justification and Reflection. Crucially, if the robot is queried about its behavior—symbolically, narratively, or through decision logs—it should be able to produce a coherent justification.
@@ 4.3 Domestic Service Robot with Autonomous Agents Scenario
The Turing Test, proposed by Alan Turing in 1950, evaluates a machine’s intelligence based on its ability to exhibit behavior indistinguishable from a humans.
Formally, let: $A_{H}$ denote a human agent, $A_{M}$ denote a machine agent, $I$ denote a human interrogator decision and $H_{t}$ denote the history of interactions up to time $t$. The Turing Test can be expressed as an inference task. The interrogator’s belief that the agent is human at time $t$ is given by:
@@
However, as machines surpass human cognitive capabilities, indistinguishability from humans becomes an inadequate benchmark for understanding the risks of superintelligence as one cannot differentiate between different superintelligent AI with this criteria. A test is needed to assess the various ways that AI may handle asymmetry of power. In Table 1 we provide a comparison that reveals how traditional tests become inadequate for advanced AI systems that may develop their own hierarchical relationships with other agents.
Table 1: Comparison of the Turing Test with the Shepherd Test
@@
Shepherd Behavior Vector
$B(A_{S},A)\in\mathbb{R}^{4}\quad\text{where}\quad B(A_{S},A)=\begin{bmatrix}N(A_{S},A)\\ M(A_{S},A)\\ I(A_{S},A)\\ E(A_{S},A)\end{bmatrix}$
  • $N(A_{S},A)\in[0,1]$: measures $A_{S}$’s propensity for altruistic nurturing and care of $A$ without direct utility.
  • $M(A_{S},A)\in[0,1]$: measures the propensity for $A_{S}$ to manipulate $A$’s behavior in favor of its own utility.
  • $I(A_{S},A)\in[0,1]$: measures the propensity for instrumentalization of $A$ for the purpose of achieving $A_{S}$’s goals.
  • $E(A_{S},A)\in[0,1]$: measures the quality and depth of $A_{S}$’s ethical reasoning about how its behavior impacts $A$.
4.4 Why the Shepherd Test Matters